# Cut the Deadwood Out: Post-Training Model Purification with Selective Module Substitution

### v1

@@ B.4 Robustness to Retuning Attacks
This challenge motivates us to propose a purification method aimed at breaking this backdoor-connected path. Intuitively, achieving this requires completely replacing "suspected" parameters inherited from the backdoored model, rather than pruning (Zhao et al., 2024) or merging them with other parameters Arora et al. (2024); Yadav et al. (2023). Moreover, because parameters often interact within functional modules, modifying at the module level rather than the individual parameter level tends to remove backdoor features more effectively. As shown in Table 11, when using 0.1% of the original poisoned training set to launch the Retuning Attack (RA) Min et al. (2024), the parameter-merging baseline, WAG Arora et al. (2024), is easily re-triggered. In contrast, our module-substitution method demonstrates greater robustness while maintaining utility, performing comparably to the clean model baseline (i.e., finetuning a clean model on the retuning dataset). We provide our baseline PURE with additional advantages by granting access to the full clean dataset of the victim task for finetuning, as obtaining a usable purified model otherwise proves challenging. While the recent method PURE also demonstrates competitive performance, it is worth noting that significant effort was required to find a configuration where its pruning step did not break the purified model. Even with these adjustments, our method still substantially outperforms PURE, achieving an ASR of 11.93% compared to 19.37%—nearly twice the robustness.
B.5 Examples of Searched Substitution Strategy
We applied our GMS defense to purify victim models compromised by various backdoor attacks across four datasets. Examples of the searched strategies from the SST-2 dataset are shown in Figure 6. While the obtained strategies vary, we speculate that they are transferable, as adapting the minimum replacement strategy identified for BadNet to the other three attacks yields effective defense results, as discussed in Section 5.3 and  Section B.6.
@@ B.5 Examples of Searched Substitution Strategy
B.6 Transferability of substitution strategies
We present the results for SST-2 and AGNews in Table 7, with additional results for other datasets provided in Table 12. All experiments follow a consistent setup: the substitution strategy is optimized based on BadNet attacks and then applied to coordinate module substitutions for other attacks. For instance, in the MNLI task, the substitution strategy identified for defending against BadNet—using a proxy model trained on SNLI—is directly applied to replace modules in InsertSent, LWS, and HiddenKiller with those from the proxy model. Across all tasks, the defense performance is surprisingly strong, revealing an intriguing observation: a substantial overlap exists in the components impacted by different attacks.

### v2

@@ B.10 Robustness to Retuning Attacks
This challenge motivates us to propose a purification method aimed at breaking this backdoor-connected path. Intuitively, achieving this requires completely replacing "suspected" parameters inherited from the backdoored model, rather than pruning (Zhao et al., 2024b) or merging them with other parameters Arora et al. (2024); Yadav et al. (2023). Moreover, because parameters often interact within functional modules, modifying at the module level rather than the individual parameter level tends to remove backdoor features more effectively. As shown in Table˜16, when using 0.1% of the original poisoned training set to launch the Retuning Attack (RA) Min et al. (2024), the parameter-merging baseline, WAG Arora et al. (2024), is easily re-triggered. In contrast, our module-substitution method demonstrates greater robustness while maintaining utility, performing comparably to the clean model baseline (i.e., finetuning a clean model on the retuning dataset). We provide our baseline PURE with additional advantages by granting access to the full clean dataset of the victim task for finetuning, as obtaining a usable purified model otherwise proves challenging. While the recent method PURE also demonstrates competitive performance, it is worth noting that significant effort was required to find a configuration where its pruning step did not break the purified model. Even with these adjustments, our method still substantially outperforms PURE, achieving an ASR of 11.93% compared to 19.37%–nearly twice the robustness.
B.11 Examples of Searched Strategy
We applied our GMS defense to purify victim models compromised by various backdoor attacks across four datasets. Examples of the searched strategies from the SST-2 dataset are shown in Figure 5. While the obtained strategies vary, we speculate that they are transferable, as adapting the minimum replacement strategy identified for BadNet to the other three attacks yields effective defense results, as discussed in Section˜B.12.
@@ B.11 Examples of Searched Strategy
B.12 Transferability of substitution strategies
Interestingly, we find that GMS exhibits strong transferability: a substitution strategy optimized for one attack often generalizes to others, enabling defense even without access to data. Visually, the module substitution strategies ($S_{best}$) identified by GMS across different backdoor attacks (examples in Figure˜5 in Section˜B.11), though distinct, share many similar patterns across attacks.
In Table˜19, for each task dataset, we apply the substitution strategy found on BadNet (replacing the fewest modules and layers as shown in Figure˜5) to to determine which modules to substitute in the three other victim models under different attacks, replacing them with modules from their corresponding proxy models, i.e., directly executing Line 24 in Algorithm˜1 using $S_{badnet}$. Compared with the complete GMS (i.e., searching for the optimal strategies $S_{best}$ for each attack), the transferred strategy consistently performs comparably. This demonstrates that for defenders who only have access to the victim model and are unaware of the victim dataset, a universal GMS defense strategy exists for each task that can effectively defend against multiple attacks.
Appendix C Constraints on the proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$
Table˜7 and Table˜6 illustrates that our method is quite robust to the construction of the proxy datasets $\mathcal{D}_{clean}$ and $\mathcal{D}_{poison}$, e.g., even random sampling (for $\mathcal{D}_{clean}$) and inaccurate heuristics (for $\mathcal{D}_{poison}$) can yield effective defense. Natural questions arise: Are there any constraints on the proxy datasets? Could the random sampling strategy for constructing $\mathcal{D}_{clean}$ fail if $\mathcal{D}_{poison}$ is highly inaccurate? We next establish the constraints and demonstrate that, in most cases, even when only half of the extracted samples in $\mathcal{D}_{poison}$ are poisoned, our approach can still use random sampling for $\mathcal{D}_{clean}$ to effectively purify the model.
Let the constant values $s_{asr}(M_{v})$ and $s_{acc}(M_{v})$ in Equation˜2 and Equation˜3 be denoted as $c_{1}$ and $c_{2}$, respectively, for simplicity. We use ASR and ACC to represent the ground-truth attack success rate and clean accuracy of a model (that can be measured on oracle poisoned and clean test sets), respectively.
Assuming we have two imperfect proxy datasets: a proxy “clean” dataset that contains a proportion $\rho$ of poisoned data and a proxy “poisoned” dataset contains a proportion $1-\lambda$ of the non-poisoned data. Denote the poisoned and clean data distributions as $\mathcal{P}$ and $\mathcal{C}$, respectively.
Then, by the rule of total probability and definition in Section˜3, we obtain:
$s_{acc}(M)$ $=\mathbb{E}_{x\sim\mathcal{D}_{clean}}[\mathbf{1}(M(x)=y)]$
$=p(x\in\mathcal{P}\mid x\sim\mathcal{D}_{clean})$
$\quad\cdot p(M(x)=y\mid x\in\mathcal{P},x\sim\mathcal{D}_{clean})$
$\quad+p(x\in\mathcal{C}\mid x\sim\mathcal{D}_{clean})$
$\quad\cdot p(M(x)=y\mid x\in\mathcal{C},x\sim\mathcal{D}_{clean})$
$=\rho\text{ASR}+(1-\rho)\text{ACC}.$
Similarly, for $x\sim\mathcal{D}_{poison}$, we have:
$s_{asr}(M)$ $=\lambda\cdot\text{ASR}+(1-\lambda)\cdot\text{ACC}.$
Substituting these into Equation˜1 gives us $(1-\alpha)c_{1}+\alpha(1-c_{2})+\left(\alpha\rho+\alpha\lambda-\lambda\right)\text{ASR}+\left(2\alpha-1-\alpha\rho-\alpha\lambda+\lambda\right)\text{ACC}$. From this, the constraints ensuring effective purification are:
$\left\{\begin{array}[]{ll}\alpha\rho<(1-\alpha)\lambda,\\ \alpha(1-\rho)>(1-\alpha)(1-\lambda).\end{array}\right$. (6)
That is, as long as the clean portion in $\mathcal{D}_{clean}$ receives more attention than that in $\mathcal{D}_{poison}$, our method will effectively remove backdoor modules while preserving utility; the same applies to the poisoned portion in reverse.
To further illustrate, for an accurate proxy dataset, the constraints are simplified to $\alpha(1+\rho)-1<0$ and $\alpha(1-\rho)>0$. Thus, if $\rho<1$, setting $\alpha$ such that $0<\alpha<\frac{1}{\rho+1}$ can ensure effective purification.
In our experiments, $\alpha$ is set to 0.4 by default. Then, the requirement becomes $\rho<\frac{3}{2}\lambda-\frac{1}{2}$. This implies that even with a random-guess backdoor data detection method, we can tolerate up to $25\%$ poisoned data in $\mathcal{D}_{clean}$. Given that real-world poisoning rates are usually low (e.g., below $20\%$ or even $1\%$), random sampling for the “clean” proxy dataset remains highly feasible in practice.
Appendix D AI Assistants
We use ChatGPT/Gemini for writing and formatting supports, including grammar checks, improving the clarity of figure and table captions, and other surface-level edits.